Two consecutive whole numbers total 17. Find the larger.
I do not understand how to do the problem , can you help ?

Answer by ankor@dixie-net.com(22740)

Two consecutive whole numbers total 17. Find the larger.
:
This is not hard
Let x = the larger number
it says,"consecutive number, therefore
(x-1) = the smaller number
:
x + (x-1) = 17
2x - 1 = 17
2x = 17 + 1
2x = 18
x = 18/2
x = 9 is the larger number, and obviously, 8 is the smaller